OpenStudy (romantic_ch3micals):

Help? Medal. and...I'll fan you The area of a the rectangular painting is given by the trinomial x^2+2x-15. What are possible dimensions? Use factoring

OpenStudy (anonymous):

x^2+2x-15=0 (ax^2+bx+c=0) (x-x1)(x+x2) D=b^2-4*a*c D=4+60=64 x1=(-b-sqrt(D))/2*x=(-2-8)/2*1=(-10)/2=-5 x2=(-b+sqrt(D))/2*x=(-2+8)/2=6/2=3 so, (x+5)(x-3)

OpenStudy (romantic_ch3micals):

ohh, darn. You could have said to use that formula. I just forgot. Hey, can you help me with one more?

OpenStudy (anonymous):

yup

OpenStudy (romantic_ch3micals):

Alright! How many number solutions does the equation have? y=3x^2-5x-5

OpenStudy (anonymous):

every square equation (ax^2+bx+c) has 2 solutions for 3x^2-5x-5 the solutions are x1=(5-sqrt(85))/6 & x2=x1=(5+sqrt(85))/6

OpenStudy (romantic_ch3micals):

ok, square equation. two solutions. Got it. Hey, do you know how do you find the solutions of a system. I want to try it out myself with this question I'm on

OpenStudy (anonymous):

x2=(5+sqrt(85))/6 a mistake (the x1) yeah, whats the question?

OpenStudy (romantic_ch3micals):

It's y=2x+2

OpenStudy (anonymous):

x1=(5-sqrt(85))/6 & x2=(5+sqrt(85))/6 <--- it should be like that (omg somthings wrong with me ) and the other equation of the system?

OpenStudy (romantic_ch3micals):

y=x^2+3x-4

OpenStudy (romantic_ch3micals):

alright, so I'm guessing it doesnt have a solution

OpenStudy (anonymous):

how so?

OpenStudy (romantic_ch3micals):

well...it doesnt match the answer choices

OpenStudy (romantic_ch3micals):

Nope. I'm trying to solve it too...

OpenStudy (anonymous):

what are the solutions?

OpenStudy (romantic_ch3micals):

1. (-3,6) and (2, -4) 2. (-3, -4) and (2, 6) 3. (-3, -4) and (-2, -2) 4. No solution

OpenStudy (anonymous):

y=y 2x+2=x^2+3x-4 x^2+3x-2x-4-2=0 x^2+x-6=0 x1=-3 x2=2 or you just misspelled some '-' or '+'...

OpenStudy (romantic_ch3micals):

oh...nevermind. It's the third choice. Alright. Hey, I'm almost done with this sample work. Sorry if I'm bothering you but I'm stuck on another problem.

OpenStudy (romantic_ch3micals):

k, i need to solve this equation. (6x^3-1x^2+1) divide (2x+1)

OpenStudy (romantic_ch3micals):

oh, no i need to find the simper form

OpenStudy (romantic_ch3micals):

simpler